# Hallucination Detection in LLMs with Topological Divergence on Attention Graphs

### v2

@@ Abstract
Hallucination, i.e., generating factually incorrect content, remains a critical challenge for large language models (LLMs). We introduce TOHA [^1] The code of the proposed method and the considered baselines is available at https://anonymous.4open.science/r/tda4hallu-BED5, a TOpology-based HAllucination detector in the RAG setting, which leverages a topological divergence metric to quantify the structural properties of graphs induced by attention matrices. Examining the topological divergence between prompt and response subgraphs reveals consistent patterns: higher divergence values in specific attention heads correlate with hallucinated outputs, independent of the dataset. Extensive experiments — including evaluation on question answering and summarization tasks — show that our approach achieves state-of-the-art or competitive results on several benchmarks while requiring minimal annotated data and computational resources. Our findings suggest that analyzing the topological structure of attention matrices can serve as an efficient and robust indicator of factual reliability in LLMs.
@@
Alexandra Bazarova1, Aleksandr Yugay1, Andrey Shulga1, Alina Ermilova1, Andrei Volodichev1, Konstantin Polev2, Julia Belikova2, Rauf Parchiev 2, Dmitry Simakov2, Maxim Savchenko2, Andrey Savchenko2, Serguei Barannikov1, 3, Alexey Zaytsev1, 1Skolkovo Institute of Science and Technology, 2Sber AI Lab, 3CNRS, Universite Paris Cite, Correspondence: a.bazarova@skoltech.ru
@@ 1 Introduction
Large language models (LLMs) have progressed significantly in recent years, finding applications in various fields Chkirbene et al. (2024). However, these models are prone to generate so-called hallucinations, i.e., content that is factually or contextually incorrect Huang et al. (2023). Detecting hallucinations is crucial for safely deploying LLMs in sensitive fields since erroneous outputs may seriously harm user trust. An effective detector would therefore expand the scope of LLM applications while mitigating risks Gao et al. (2024).
Multiple methods address this problem Huang et al. (2023); Sahoo et al. (2024), though many face significant practical constraints. A common limitation is the reliance on large annotated datasets Sky et al. (2024); Azaria and Mitchell (2023); Chuang et al. (2024), which are rarely available publicly Zhang et al. (2023) and require extensive annotation effort for each new model released. Another popular approach depends on generating multiple samples for scoring Manakul et al. (2024); Chen et al. (2024); Farquhar et al. (2024), which increases computational costs substantially.
Figure 1: Difference between average topological divergence values for hallucinated and grounded samples per attention head/layer, evaluated on MS MARCO and CoQA datasets. A lighter color corresponds to a greater difference. Green frames highlight the heads that separate samples best. The same attention heads assign greater divergence values to the hallucinated samples in both datasets. Model: Mistral-7B-Instruct-v0.1.
We address these challenges by introducing TOHA (TOpology-based HAllucination detector), a training-free method for the retrieval-augmented generation (RAG) setting Gao et al. (2023). Following the prior work Du et al. (2024), TOHA requires minimal annotated data (just $50$ annotated samples suffice for reliable detection, see Figure 4) while avoiding the computational overhead of multiple generations, making it both data- and compute-efficient.
The core insight behind TOHA is that hallucinated answers tend to have weaker connections to the context than grounded ones in the RAG setting. TOHA formalizes this by analyzing attention graphs — complete graphs induced by LLM attention maps, a representation previously used in other NLP tasks Kushnareva et al. (2021); Tulchinskii et al. (2023). Unlike prior works that rely on simplistic classifiers over basic graph properties Proskurina et al. (2023a); Cherniavskii et al. (2022), TOHA advances the approach by computing the topological dissimilarity between the attention subgraphs of the model’s response and the given context. This dissimilarity, measured via our adaptation of Manifold Topology Divergence  Barannikov et al. (2021) for the graph setting, quantifies the “strength” of context-response ties: higher values indicate weaker links and thus likely hallucinations. We prove key stability properties for this metric, ensuring its reliability as a hallucination score.
@@
Figure 2: a) An attention map. Blue and green denotes the prompt and response tokens, respectively. b) The corresponding attention graph $G$. Prompt tokens $P$ are located on the left, response tokens $R$ — on the right. To keep figure neat, we only plot the edges with an attention score of no less than $0.15$. c) The minimum spanning forest attaching $R$ to $P$.
Through analysis of divergence patterns across different heads, we identified a subset of attention heads that consistently assign higher divergence scores to hallucinated samples (see Figure 1), revealing their implicit “awareness” of hallucinations. TOHA utilizes the average divergence values from these specific heads as hallucination scores. Crucially, these heads exhibit consistent behavior across different datasets, enabling strong transferability of our method.
Our main contributions can be summarized the following:
@@
  • We propose TOHA, a training-free method based on the topological divergences of attention graphs. While efficient — TOHA operates up to an order of magnitude faster than methods of comparable quality and requires minimal annotated data — our method demonstrates strong in-domain performance and maintains domain transferability across different tasks.
  • The existence of hallucination-aware attention heads is discovered: calculating topological divergences from just six specific heads is enough for reliable hallucination detection, irrespective of the dataset.
@@
  • Our experiments show TOHA consistently matches or exceeds state-of-the-art performance on all benchmarks when applied to modern open-source LLMs of varying scales (7B to 13B parameters).
@@ 2 Background
2.1 Attention matrix as a weighted graph
@@
where $W(Q,K)$ is an attention matrix
$\operatorname{W}=\operatorname{softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right),$ (2)
and each entry $w_{ij}$ in it captures how strongly token $i$ attends to token $j$, $i\geq j$ for a decoder, with larger $w_{ij}$ indicating closer relationship.
An attention matrix $W$ can be reframed as a complete weighted graph $G$ where tokens are vertices and weights $w_{ij}$ represent the strength of connections between them. From the perspective of topological data analysis, however, it is more convenient to consider these weights as pseudo-distances rather than correlation measures. Hence, we reassign the edge weights of such a graph to equal $1-w_{ij}$, creating what we call attention graphs.
In the generation process, the vertices of such graphs naturally partition into two distinct subsets: prompt tokens $P$ and response tokens $R$ generated by the model (see Figure 2b). This split allows us to analyze the topological relationships between input and output content.
@@ 2.2 Manifold Topology Divergence
One way to compare two data manifolds, $\mathcal{M}$ and $\mathcal{N}$, approximated by point clouds $M$ and $N$ is the $\operatorname{MTop-Div}(M,N)$ topological measure Barannikov et al. (2021). This divergence is based on the $\operatorname{Cross-Barcode}(M,N)$, which is a set of intervals $\{(b_{i},d_{i})\}_{i=1}^{n}$ corresponding to “births” and “deaths” of independent topological features that distinguish the point cloud $N$ from the union $M\cup N$. The further $\operatorname{Cross-Barcode}(M,N)$ is from an empty set, the more the data manifold $\mathcal{N}$ differs from $\mathcal{M}$ in its topological structure. To measure the distance from an empty set, the sum of lengths of intervals in $\operatorname{Cross-Barcode}(M,N)$ was taken in loc. cit.:
$\operatorname{MTop-Div}(M,N)=\sum\limits_{i=1}^{n}|d_{i}-b_{i}|.$
@@ 3 Method
Given an attention matrix for the (prompt + response) text, we construct the attention graph, imitating a data manifold of the text, and study its relation with the weighted subgraph, imitating the data submanifold of the prompt. We measure the topological divergence between these graphs, assuming responses that are consistent with the context would result in lower divergence values.
3.1 $\operatorname{MTop-Div}$ for attention graphs
The $\operatorname{MTop-Divergence}(M,N)$ measure was originally developed for data manifolds, relying on metric space properties to some extent. We adapt this concept to quantify divergence between attention graphs and their subgraphs, where traditional metric axioms do not apply, while preserving the measure’s comparative utility.
Let $R$ and $P$ be the sets of response and the prompt vertices in the attention graph $G$. We set to zero the edge weights between the $P$ vertices, denote $w_{(R\cup P)/P}$ the resulting matrix of edge weights, and define $\operatorname{Cross-Barcode}_{i}(R,P)$ as the $i$-th homology barcode of the Vietoris-Rips simplicial complex $VR_{\alpha}(G,w_{(R\cup P)/P})$. We define $\operatorname{MTop-Div}(R,P)$ as the total sum of interval lengths in $\operatorname{Cross-Barcode}_{0}(R,P)$, where we consider the 00-dimensional homology group $H_{0}$. In this context, we can prove that this score is equivalent to the total edge length of the minimum spanning forest (MSF) connecting $R$ to $P$.
Properties of MTop-Div for attention graphs.
Here, we only list the properties relevant to hallucination detection; for more properties and proofs, see Appendix A.
@@ 3.1 $\operatorname{MTop-Div}$ for attention graphs
1. (Formula.) $\operatorname{MTop-Div}(R,P)$ value equals the length of the MSF attaching $R$ to $P$.
2. (Stability.) If the weights of $G$ change by no more than $\varepsilon$, then the corresponding $\operatorname{MTop-Div}(R,P)$ changes by no more than $\delta=\varepsilon|R|$.
3. (Connection with hallucinations.) The normalized divergence value $\frac{1}{|R|}\operatorname{MTop-Div}(R,P)=0$ iff the MSF attaches every response token to a prompt token by a subtree with attention weights $=1$.
The stability property guarantees that similar attention patterns yield similar hallucination scores, making the metric’s behavior consistent and predictable. The latter captures the intuitive relationship between divergence and response quality: well-grounded responses (closely tied to the context) produce small divergence values, while hallucinations (occurring when evidence is missing) lead to smaller attention weights and consequently larger divergence values. Together, these properties enable the metric to reliably measure how strongly a response connects to its context while remaining robust to minor attention variations.
@@ 3.2 Hallucination-aware heads
We hypothesize, inspired by prior investigations in LLM interpretability Voita et al. (2019); Gould et al. (2024), that particular attention heads exhibit distinct patterns related to hallucinations. To identify such heads, we analyzed head-specific topological divergences as follows.
@@
$\Delta_{ij}=\frac{1}{|S_{\mathrm{hallu}}|}\sum\limits_{s\in S_{\mathrm{hallu}}}d_{ij}(s)-\frac{1}{|S_{\mathrm{gr}}|}\sum\limits_{s\in S_{\mathrm{gr}}}d_{ij}(s),$
@@
$d_{ij}(s)=\frac{1}{|R_{ij}^{s}|}\operatorname{MTop-Div}(R^{s}_{ij},P^{s}_{ij}).$
The sample differences obtained for three datasets are displayed in Figure 3. Each dot represents an individual attention head, with its $x$-coordinate indicating its $\Delta_{ij}$ value on dataset (A) and its $y$-coordinate the corresponding value on dataset (B). For each dataset pair, we highlight the top three most separating heads for dataset (A) in pink. Notably, these heads consistently appear in the upper-right corner of the plot, indicating that they also exhibit strong separation on dataset (B). This observation suggests that these heads are inherently attuned to hallucination patterns, regardless of the dataset.
Figure 3: $\Delta_{ij}$ values for $ij$-th heads. Vertical axis corresponds to the difference on dataset (B), horizontal — to the one on dataset (A). The heads that separate samples best are highlighted in pink. Model names for a row are on the left side, datasets: MS MARCO, CNN/DM + Recent News, CoQA.
@@ 3.3 TOHA
The existence of universal hallucination patterns in the attention heads underlies our efficient method TOHA, detailed in Algorithm 1. It uses two small, annotated probe sets, $S_{h}$ (containing hallucinated samples) and $S_{g}$ (containing grounded samples), to rank model heads by their separation capability based on their $\Delta_{ij}$ values, where $ij$ denotes a head index, and select the most relevant ones. In our experiments, the combined size of the probe sets is kept small (see Figure 4 for the number of the required samples analysis). During testing, hallucination scores are computed as the average topological divergence from the top $N_{\mathrm{opt}}$ heads, where $N_{\mathrm{opt}}$ is a hyperparameter tuned on the set $V=S_{h}\cup S_{g}$. For computational efficiency, we limit $N_{\mathrm{opt}}$ to a maximum of 6 in all experiments.
@@
$d_{ij}(s)$ — topological divergences for samples;
$S_{h},\,S_{g}$ — probe sets;
$V=S_{h}\cup S_{g}$ — validation set;
$T$ — test set;
$N_{\mathrm{max}}$ — max number of selected heads.
procedure TOHA heads selection
     $\Delta_{ij}\leftarrow\frac{1}{|S_{h}|}\sum\limits_{s\in S_{h}}d_{ij}(s)-\frac{1}{|S_{g}|}\sum\limits_{s\in S_{g}}d_{ij}(s)$
     $H\leftarrow\mathrm{sort\_descending}(h_{ij},\text{key}=\Delta_{ij})$
     $N,N_{\mathrm{opt}}\leftarrow 1,1$
     $H_{\mathrm{subset}}\leftarrow\varnothing$triangleright Optimal heads set.
     $\text{AUROC}_{\max}\leftarrow 0$
     $p_{s}=0,s\in V$triangleright Hallucination scores.
     while $N\leq N_{\max}$ do
         $H_{\mathrm{subset}}\leftarrow H_{\mathrm{subset}}\cup\{h_{N}\}$
         for $s\in V$ do
              $p_{s}\leftarrow\frac{N-1}{N}p_{s}+\frac{1}{N}d_{h_{N}}(s)$
         end for
         $\text{auroc}\leftarrow\text{AUROC}(\{y_{s}\}_{s\in V},\{p_{s}\}_{s\in V})$
         if $\text{auroc}>\text{AUROC}_{\max}$ then
              $\text{AUROC}_{\max}\leftarrow\text{auroc}$
              $N_{\mathrm{opt}}\leftarrow N$
         end if
         $N\leftarrow N+1$
     end while
end procedure
procedure TOHA prediction
     for $s\in T$ do $\triangleright$ Prediction on the test set.
         $p_{s}\leftarrow\frac{1}{N_{\mathrm{opt}}}\sum\limits_{i=1}^{N_{\mathrm{opt}}}d_{h_{i}}(s)$
     end for
end procedure
4 Experiments
Datasets.
We evaluated our approach on four datasets: RAGTruth Niu et al. (2023) (we considered its’ two separate benchmarks: QA based on MS MARCO Nguyen et al. (2016) and summarization based on CNN/DM Nallapati et al. (2016)) combined with news articles from an unnamed news platform), CoQA Reddy et al. (2019), SQuAD Rajpurkar et al. (2016), and XSum Narayan et al. (2018). The RAGTruth dataset consists of manually annotated responses of several LLMs in the RAG setting. The annotations are word-level; we, in turn, predict response-level labels, considering a response hallucinated if it contains at least one hallucination span. For the three latter datasets, we sampled LLM responses and annotated them automatically with GPT-4o Hurst et al. (2024). Consistency with Human-GPT-4o label (Table 7) validated this approach, matching prior work Bavaresco et al. (2024); see Appendix C for more details.
Models.
We used five popular open-source LLMs: LLaMA-2-7B-chat, LLaMA-2-13B-chat, LLaMA-3.1-8B-Instruct, Mistral-7B-Instruct-v0.1, and Qwen2.5-7B-Instruct. Note that the RAGTruth dataset does not contain responses for LLaMA-3.1-8B and Qwen-2.5-7B; therefore, we only conducted experiments on SQuAD, CoQA, and XSum for these models.

### v3

@@ Abstract
Hallucination, i.e., generating factually incorrect content, remains a critical challenge for large language models (LLMs). We introduce TOHA [^1] The code of the proposed method and the considered baselines is available at https://anonymous.4open.science/r/tda4hallucinations-C449, a TOpology-based HAllucination detector designed for RAG scenarios, which leverages a topological divergence metric to quantify the structural properties of graphs induced by attention maps. Examining the topological divergence between prompt and response subgraphs reveals consistent patterns: higher divergence values in specific attention heads correlate with hallucinated outputs, independent of the dataset. Extensive experiments, including evaluation on question answering and summarization tasks, show that our approach achieves state-of-the-art or competitive results on several benchmarks while requiring minimal annotated data and computational resources. Our findings suggest that analyzing the topological structure of attention matrices can serve as an efficient and robust indicator of factual reliability in LLMs.
@@
Alexandra Bazarova1, Aleksandr Yugay1, Andrey Shulga1, Alina Ermilova1, Andrei Volodichev1, Konstantin Polev2, Julia Belikova2, Rauf Parchiev 2, Dmitry Simakov2, Maxim Savchenko2, Andrey Savchenko2, Serguei Barannikov1, 3, Alexey Zaytsev1, 1Applied AI Institute, 2SB AI Lab, 3CNRS, Universite Paris Cite, Correspondence: bazarovaai.239@gmail.com
@@ 1 Introduction
Large language models (LLMs) have progressed significantly in recent years, finding applications in various fields Chkirbene et al. (2024). To ensure factual reliability, modern LLMs are often combined with retrieval-augmented generation (RAG) technique, which integrates relevant external knowledge from diverse databases directly into the generation process Lewis et al. (2020).
@@
Despite these improvements, LLMs remain prone to producing so-called hallucinations, i.e., content that is factually or contextually incorrect Huang et al. (2023); Li et al. (2024). Detecting hallucinations is crucial for safe deployment of LLMs in sensitive fields since erroneous outputs may seriously harm user trust. An effective detector would therefore expand the scope of LLM applications while mitigating risks Gao et al. (2024).
Multiple methods have been proposed to tackle this problem Sahoo et al. (2024); Shorinwa et al. (2025). However, they often face significant practical constraints, such as the scarcity of annotated datasets Zhang et al. (2023) required for supervised methods Sky et al. (2024); Orgad et al. (2025), the high computational cost of generating multiple additional samples Chen et al. (2024); Hou et al. (2025), or the inability of LLMs’ output probabilities to fully represent the model’s true uncertainty Fadeeva et al. (2024); Shelmanov et al. (2025).
These challenges can be addressed by utilizing the inner states of LLMs, which are informative for the hallucination detection problem Azaria and Mitchell (2023); Sriramanan et al. (2024); Gekhman et al. (2025). We introduce TOHA (a TOpology-based HAllucination detector), a training-free method designed for RAG scenario that leverages the structure of LLM attention maps to identify hallucinations. Our method requires minimal annotated data while avoiding the computational overhead of multiple generations, which makes TOHA both data- and compute-efficient.
The core insight behind TOHA is that the structure of prompt-response interconnections within an LLM’s attention mechanism should be indicative of hallucination presence. TOHA formalizes this idea by analyzing attention graphs — complete weighted graphs derived from LLM attention maps, the representation prior used for topological data analysis (TDA) in NLP Kushnareva et al. (2021); Tulchinskii et al. (2023). Unlike the existing attention-based methods, which either treat all attention heads as equally important Sriramanan et al. (2024); Binkowski et al. (2025) or ignore the geometric structure of attention maps Sun et al. (2025); Vazhentsev et al. (2025), our method leverages a small set of attention heads selected by a topological criterion designed to identify hallucinations.
Specifically, TOHA is based on the $\operatorname{MTop-Div}_{G}(R,P)$, our adaptation of Manifold Topology Divergence  Barannikov et al. (2021) for the graph setting, which quantifies the dissimilarity between the prompt ($P$) and response ($R$) token sets in the attention graph. We demonstrate that this score not only reflects the geometry but also measures informational novelty of the response in relation to the prompt, thus being well-suited for hallucination detection in RAG scenarios.
@@
Figure 1: a) An attention map. Blue and green denotes the prompt and response tokens, respectively. b) The corresponding attention graph $G$. Prompt tokens $P$ are located on the left, response tokens $R$ — on the right. To keep figure neat, we only plot the edges with an attention score of no less than $0.15$. c) The minimum spanning forest attaching $R$ to $P$ and the corresponding $\operatorname{MTop-Div}$ value.
Through analysis of divergence values across attention heads, we identified a subset of heads that consistently yield greater scores for hallucinated samples (see Figure 2), irrespective of the dataset. TOHA’s final hallucination score is the average $\operatorname{MTop-Div}_{G}(R,P)$ value from these “hallucination-aware” heads. We discover that some of these heads are associated with copying behavior, which is aligned with prior findings Sun et al. (2025).
Our main contributions can be summarized as follows:
@@
  • We introduce TOHA, a training-free framework for hallucination detection that leverages the topology of LLM attention graphs. Our method operates up to an order of magnitude faster than methods of comparable quality and requires minimal annotated data.
  • Central to TOHA is $\operatorname{MTop-Div}_{G}(R,P)$, an adaptation of Manifold Topology Divergence for graphs, which quantifies the topological dissimilarity between prompt and response sets of tokens. We demonstrate that this score characterizes the novelty of the response with respect to the prompt, making it highly effective for detecting hallucinations in RAG systems.
  • By analyzing the proposed score, we discover the existence of “hallucination-aware” attention heads, which consistently yield greater divergence values for hallucinated samples across different datasets. This finding ensures TOHA’s efficiency and solid cross-domain transferability: averaging $\operatorname{MTop-Div}_{G}(R,P)$ for just ten heads is enough for robust hallucination detection.
@@
  • Our experiments show that TOHA consistently matches or exceeds state-of-the-art performance across multiple benchmarks when applied to open-source LLMs of varying scales, including 7B and 13B parameters.
@@ 2 Background
2.1 Attention Map as a Weighted Graph
@@
where $W(Q,K)$ is an attention map
$\operatorname{W}(Q,K)=\operatorname{softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right),$ (2)
and each entry $w_{ij}=W_{ij}(Q,K)$ captures how strongly token $i$ attends to token $j$, with greater values indicating closer relationship. We consider decoder-only LLMs, for which the attention maps are lower-triangular.
An attention map can be reframed as a complete undirected weight graph $G$ with edge weights $1-w_{ij}$ to represent pseudo-distances between tokens. We call $G$ an attention graph. It naturally partitions into prompt ($P$) and response ($R$) tokens (see Figure 1b). We analyze the topological relationships between these node subsets, which, as we assume, should be indicative of hallucinations in the RAG scenario.
@@ 2.2 Manifold Topology Divergence
The Manifold Topology Divergence, or $\operatorname{MTop-Div}$, was proposed in Barannikov et al. (2021). $\operatorname{MTop-Div}(M,N)$ quantifies the difference between two data manifolds $\mathcal{M}$ and $\mathcal{N}$, approximated by point clouds $M$ and $N$. It is computed as the sum of interval lengths in the $\operatorname{Cross-Barcode}(M,N)$, a set of intervals representing topological features distinguishing $N$ from $M\cup N$. A larger divergence indicates a greater topological difference between the manifolds. For more details on topological data analysis, see Appendix C.
@@ 3 Method
This section introduces the $\operatorname{MTop-Div}_{G}(R,P)$ score, which is designed to quantify the topological divergence between the prompt and response subgraphs in attention maps. We demonstrate that not only can it be interpreted as a geometric characteristic of the attention graph, but also as an information-theoretic measure of the novelty of the response in relation to the prompt. With $\operatorname{MTop-Div}_{G}(R,P)$, we identify “hallucination-aware” heads that consistently separate hallucinated samples from grounded, irrespective of dataset. This finding allows us to formulate the TOHA algorithm (Algorithm 1), which computes a final hallucination score by averaging the divergence values from these specific heads.
3.1 $\operatorname{MTop-Div}$ for Attention Graphs: Definition
Let $R$ and $P$ be the response and prompt vertex sets in an attention graph $G$. After zeroing edge weights between the $P$ nodes, we compute the 0-th order homology barcode $\mathcal{B}_{0}$ of the Vietoris-Rips simplicial complex of the modified graph. Essentially, this barcode tracks the evolution of connected components as we threshold the graph edges (see Appendix C for details).
Our proposed topological divergence is a sum of lengths of $\mathcal{B}_{0}$ intervals:
$\operatorname{MTop-Div}_{G}(R,P)=\sum\limits_{\scriptstyle[b_{i},d_{i}]\in\mathcal{B}_{0}}|d_{i}-b_{i}|.$
This score can be interpreted from two perspectives: geometric (as the length of a minimal spanning forest in the attention graph) and information-theoretic (as a measure of the response novelty in the space induced by query-key matrices). We explore these two interpretations in the following subsections.
3.2 $\operatorname{MTop-Div}_{G}(R,P)$ and Geometry of Attention Graph
Formally, we can prove the following property (see proof in Appendix C):
@@ 3.2 $\operatorname{MTop-Div}_{G}(R,P)$ and Geometry of Attention Graph
Consider an attention graph $G$ with vertex set $V_{G}$ and its complementary vertex subsets $P,R$, where $P\cup R=V_{G}$ and $P\cap R=\varnothing$. $\operatorname{MTop-Div}(R,P)$ value equals the length of the minimal spanning forest (MSF) attaching $R$ to $P$.
This proposition’s geometric meaning is illustrated in Figure 1. While the connections between prompt nodes (Figure 1b) are semantically and syntactically meaningful, we hypothesize they primarily introduce noise for hallucination detection (see Section 4.4 for the corresponding experiment). After setting these distances to zero, we construct an MSF on the modified graph (Figure 1c). The sum of the lengths of the remaining edges in this MSF is precisely the value of $\operatorname{MTop-Div}_{G}(R,P)$.
3.3 $\operatorname{MTop-Div}_{G}(R,P)$ and Perplexity in the Query-Key Space
Consider an attention graph $G$. It can be interpreted as a non-metric***The $Q^{T}K$ scalar product from self-attention cannot be formally interpreted as a metric: for one, it does not adhere to the triangle’s inequality. space, with “distances” induced by $Q$ and $K$ matrices from the attention mechanism.
In order to quantify some notion of entropy in this non-metric space, the most intuitive way is to measure the MST length Müller et al. (2012) as a proxy for token dispersion. More precisely, the entropy estimate given the MST of length $L$ on points $\mathcal{X}$ is the following:
$H_{MST}(\mathcal{X})=d\log L-(d-1)\log n+\log\beta_{d},$
where $|\mathcal{X}|=n$, $d$ is the intrinsic dimensionality of the data, and $\beta_{d}$ is some data-independent constant.
We can show that
$\operatorname{MTop-Div}_{G}(R,P)\geq L_{MST}(R\cup P)-L_{MST}(P),$
where the first term corresponds to the MST length of the complete attention graph $G$, and the second to the length of the MST spanning the prompt tokens $P$. Therefore, our score has the following lower bound:
$\operatorname{MTop-Div}_{G}(R,P)\geq\alpha 2^{H(P\cup R)}-\beta 2^{H(P)},$
$\alpha,\beta>0$, or, in terms of the perplexity function,
$\operatorname{MTop-Div}_{G}(R,P)\geq\alpha\mathrm{PPL(P\cup R)}-\beta\mathrm{PPL(R)}.$
Perplexity measures how “surprising” a piece of data is. Therefore, the difference between the perplexity values quantifies the novelty of the response with respect to the prompt: greater values suggest that new information is present, which, in the RAG scenario, is considered to be a hallucination. Thus, our $\operatorname{MTop-Div}_{G}(R,P)$ should be an effective statistic for identifying hallucinated responses.
@@ 3.4 Hallucination-Aware Heads
$\Delta_{ij}=\frac{1}{|S_{\mathrm{hallu}}|}\sum\limits_{s\in S_{\mathrm{hallu}}}d_{ij}(s)-\frac{1}{|S_{\mathrm{gr}}|}\sum\limits_{s\in S_{\mathrm{gr}}}d_{ij}(s),$ (3)
@@
$d_{ij}(s)=\frac{1}{|R_{ij}^{s}|}\operatorname{MTop-Div}_{G^{s}_{ij}}(R^{s}_{ij},P^{s}_{ij}).$
Figure 2 displays sample differences $\Delta_{ij}$ across three datasets, with each marker representing some attention head. We discovered that the same four (for Mistral-7B) and three (for Llama-2-7B) heads, highlighted in pink, demonstrate similar behaviour across the datasets: they consistently appear in the upper-right corner, indicating strong separation between hallucinated and grounded samples, irrespective of the dataset. This finding suggests that there are specific attention heads somehow “aware” of the presence of hallucination, which is captured by the proposed $\operatorname{MTop-Div}_{G}(R,P)$.
Figure 2: $\Delta_{ij}$ values for ijij-th heads. Vertical axis corresponds to the difference on dataset (B), horizontal — to the one on dataset (A). The heads that separate samples best are highlighted in pink; their (layer, head) positions are reflected in the legend.
@@ 3.5 TOHA
The existence of “hallucination-aware” attention heads underlies our method, which is detailed in Algorithm 1. TOHA uses two annotated probe sets ($S_{h}$ for hallucinations, $S_{g}$ for grounded samples) to rank attention heads by their separation capability $\Delta_{ij}$ and select the top $N_{\mathrm{opt}}$ ones with the combined probe size $|S_{h}\cup S_{g}|$ kept small (see Figure 3). During testing, the final hallucination score for test samples $T$ equals the average topological divergence from the selected $N_{\mathrm{opt}}$ heads. For computational efficiency, we limit $N_{\mathrm{opt}}$ to a maximum of $N_{\max}=10$ in all experiments.
@@
[…]
2:Hallucination scores $\{p_{s}\}$ for $s\in T$
3:
4:function HeadsSelection
5:  for each head $h_{ij}$ do
6:   Calculate $\Delta_{ij}$ according to eq. (3)
7:  end for
8:  $H\leftarrow$ sort all heads by $\Delta_{ij}$ (descending)
9:  $N_{\text{opt}},\text{AUROC}_{\max}\leftarrow 1,0$
10:  Initialize $p_{s}\leftarrow 0$ for all $s\in V$
11:  for $N=1$ to $N_{\max}$ do
12:   for each $s\in V$ do
13:     $p_{s}\leftarrow\frac{N-1}{N}p_{s}+\frac{1}{N}d_{h_{N}}(s)$
14:   end for
15:   $\text{auroc}\leftarrow\text{AUROC}(\{y_{s}\}_{s\in V},\{p_{s}\}_{s\in V})$
16:   if $\text{auroc}>\text{AUROC}_{\max}$ then
17:     $\text{AUROC}_{\max}\leftarrow\text{auroc};\ N_{\text{opt}}\leftarrow N$
18:   end if
19:  end for
20:  return $N_{\text{opt}}$
21:end function
22:
23:function Prediction($N_{\text{opt}}$)
24:  for each $s\in T$ do
25:   $p_{s}\leftarrow\frac{1}{N_{\text{opt}}}\sum_{i=1}^{N_{\text{opt}}}d_{h_{i}}(s)$
26:  end for
27:  return $\{p_{s}\}$
28:end function
4 Results
4.1 Experiment Setting
Datasets and models.
We consider five tasks that evaluate question answering and summarization abilities of LLMs: two benchmarks of RAGTruth Niu et al. (2023) — long-form QA dataset MS MARCO and summarization dataset CNN/DM; conversational QA dataset CoQA Reddy et al. (2019), reading comprehension dataset SQuAD Rajpurkar et al. (2016), and extreme summarization dataset XSum Narayan et al. (2018). For more details, see Appendix E.
We employ five popularly adopted open-source LLMs with accessible inner states: LLaMA-2-7B-chat, LLaMA-2-13B-chat, LLaMA-3.1-8B-Instruct, Mistral-7B-Instruct-v0.1, and Qwen2.5-7B-Instruct. As the RAGTruth dataset does not contain responses for LLaMA-3.1-8B and Qwen-2.5-7B, we conducted experiments on SQuAD, CoQA, and XSum for these models.